1
Bubbas Western Wear is a western hat retailer in Lubbock, Texas. Although Bubbas carries numerous styles of western hats, each hat has approximately the same price and invoice (purchase) cost, as shown in the following table. Sales personnel receive a commission to encourage them to be more aggressive in their sales efforts. Currently, the Lubbock economy is really humming, and sales growth at Bubbas has been great The business is very competitive, however, and Bubba, the owner, has relied on his knowledgeable and courteous staff to attract and retain customers who otherwise might go to other western wear stores. Because of the rapid growth in sales, Bubba is also finding the management of certain aspects of the business more difficult, such as restocking inventory and hiring and training new salespeople. $ 80.00 Sales price Per unit variable expenses Purchase cost Sales commissions 43.50 11.50 $ 55.00 Total per unit variable costs Total annual fixed expenses Advertising Rent Salaries $ 98,50e 146,500 255,009 $500,000 Total fixed expenses
media%2F97c%2F97c34307-1050-46ee-832c-ea
media%2F196%2F19695d78-cd79-43b9-a27c-5d
media%2F390%2F3906a573-386a-4c3c-8e83-21
media%2Fc0c%2Fc0c08218-13de-4979-961a-46
Answer #1

hi, please find below the answer let me know if you need any clarification -

Given that -
Sales price 80
Variable cost
Purchase cost 43.5
Sales commmission 11.5
Total variable cost 55
Fixed cost
Advertisement 98500
Rent 146500
Salaries 255000
Total fixed cost 500000
Answer 1)
i Sales price $   80.00
ii Total variable cost $   55.00
iii =i-ii Contribution margin $   25.00
iv=iii/i Contribution margin ratio 31.25%
Break even unit Break even unit = Fixed cost / Contribution margin
=500000/25
20000 unit
Break even $ sales Break even $ sales = Fixed cost / Contribution margin ratio
=500000/31.25%
$                               1,600,000
Answer 2)
Before tax income with 22000 unit of hats = Contribution margin per unit * 22000-fixed cost
=25*22000-500000
50000
Contribution income statement
Sales =22000*80 1760000
Variable cost = 22000*55 1210000
Contribution margin 550000
Fixed cost 500000
Before tax income 50000
Answer 3)
Margin of safety $ = Actual sales - break even sales
=32000*80-20000*80
960000
Margin of safety unit = Actual sales - break even sales
=32000-20000
12000
Margin of safety ratio = (Actual sales - break even sales)/Actuals sales
=(32000-20000)/32000
37.50%
Answer 4)
Break even point with new plan
Sales per unit = 80
revised variable cost = 43.5
Contribution per unit 36.5
Fixed cost 657000
=500000+157000
Break even point = 657000/36.5 18000
Profit and loss with 22000 unit sold
Contribution = 22000*36.5 803000
Less fixed cost 657000
Before tax earning 146000